Long Title

First Author Second Author and Third Author

Abstract

Our premise is ...

Index Terms

keyword 1, keyword 2 ...

I. Introduction

USE the PARstart macro only for the first paragraph in the entire paper...

Theorem 1 (Theorem name) Consider the system ...

Proof: The proof is trivial.

Gregory L. Plett A bio with no face...

filename (missing from disk!)

Department of Electrical Engineering Some University, Somewhere CA, 90210, USA Phone: (xxx) xxx-xxxx, email: xxx@xxxx.xxx.xxx

Department of Electrical Engineering... ...

March 30, 2006—6 : 09 pm DRAFT